It has been a busy summer for the teams preparing the restart of the SPS complex. During LS1 a vast programme of repair, consolidation and improvements was completed, and surprises came up too, leading for instance for a last minute replacement of the high energy beam dump, just weeks before the restart of the machine.
One of the key elements of the consolidation programme was the work in the TDC2/TTC2 areas: the transfer lines from the SPS to the target area, and the target area itself. The planned work was completed over the summer, thus allowing operating the fixed target cycles at nominal intensity and an early restart in 2015.
Early October the proton beams were extracted to the North Area, and a lively programme of fundamental research, R&D studies for future detectors, and calibration and machine studies resumed after the long shutdown at the expected date.
The test-beams were the first to receive beam on October 6th: CMS started with a tracker and calorimeter R&D programme, while in the neighbouring beam line, the LHCf collaboration is validating their two calorimeters, prior to their installation this month in the LHC tunnel. The HSE-RP unit made use of the CERF facility on the H6 beam line, to conduct a campaign of inter-calibration of the dosimeters used at different organizations around the world, while in the H8 beam line, LHCb,
TOTEM, UA9 and various groups working on pixel detectors collaborate to make the best use of the available beams.
A few days later the beams were set up for fixed-target experiments. The NA62 experiment has now observed its first kaons with the detector close to completion, while COMPASS is preparing its Drell-Yan measurement campaign. Last week the NA61 experiment started its calibration run, in preparation of the Ar ion run scheduled early 2015. The fixed target run of 2014, short but intense, had a very promising start.